OpenStudy (anonymous):

HELP!!!!!!!!!!!!! Solve the system of equations. 3x + 3y = –6 –x + y = 0

OpenStudy (anonymous):

ponit

OpenStudy (anonymous):

Solve by elimination

OpenStudy (hockeychick23):

3x+3y= -6> 3x+3y=-6 3(-x+y=0) > -3x+3y=0 Add them> 6y= -6> y= -1 Plug y into the equation -x-1=0 x=-1

OpenStudy (anonymous):

yeeeeeeessssssssssss

OpenStudy (anonymous):

Dont give the answer

OpenStudy (anonymous):

I don't understand how to solve it.

OpenStudy (anonymous):

Hockeychick gave you the answer...

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

I know but I want to know how to solve it.

OpenStudy (hockeychick23):

oh i tried writing the steps, you have to use elimination which means you need to get either x or y to have the same number in front of it in the equation, so i multiplied the second equation by 3 so you could eliminate x

OpenStudy (anonymous):

Why 3?

OpenStudy (anonymous):

@hockeychick23

OpenStudy (hockeychick23):

because you want to eliminate either x or y in the equation, so in order to eliminate x both equations need to have the same coefficient in front of them (in this case it's 3)

OpenStudy (anonymous):

Okay so when you plug them in what do you add? You lost me on that part lol

OpenStudy (anonymous):

@hockeychick23

OpenStudy (hockeychick23):

so 3x+3y= -6 and -3x+3y=0, so add like terms, 3x+ -3x which will give you 0 and 3y+3y will give you 6y

OpenStudy (anonymous):

ohh no I get that part okay, once you get that then what?

OpenStudy (anonymous):

@hockeychick23

OpenStudy (hockeychick23):

ok so now you have 0x+6y=-6

OpenStudy (hockeychick23):

so the 0x doesn't matter because anything times 0 is 0, so now you have 6y= -6